shyshy:

what is (6 1/3 + 4 4/9)- (3 2/3 + 2 2/9)=

supie:

I thought it would be \[61/3+44/9−(32/3+22/9)=\] \[227/9−(32/3+22/9)=\] \[=227/9−118/9=?\] I'm not 100% sure tho.

R3al:

OK, but you know you can search it up, but it isn't always right tho, but thanks for trying!

supie:

I didn't search it up.